OpenStudy (anonymous):

Solve the following system of equations for the unknown variables. a – 6 = 2b a + b = 15 im suppose to be solving by substitution but i get half way through and cant figure it out?

OpenStudy (anonymous):

rewrite : a-2b=6

OpenStudy (anonymous):

u need solve a and b ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

time (-1) u get -a-b=-15

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

a-2b=6 -a-b =-15 ------------------- add -3b =-9 b=-9/-3 b =3 sutitution b = 3 solve for a a – 6 = 2(3) a - 6 =6 a =6+6 a =12 if I deverse a medal tha give me one TY

OpenStudy (anonymous):

k

OpenStudy (anonymous):

ok?